OpenStudy (anonymous):

2. The cost of a jet ski lesson on Madeira Beach can be calculated using the equation C=40h + 100, where c represents the cost of the lesson and h represents the total amount of hours of the lesson. What is the cost of a five-hour jet ski lesson? A. $140 C. $300 B. $200 D. $505

OpenStudy (anonymous):

plug 5 for h and solve

OpenStudy (anonymous):

c = 40(5) +100 = 200 +100 c =300